OpenStudy (ambermarie151):

Which of the following is a geometric sequence ? A.6,18,54,162 B.3,6,9,12,15 C.3,5,8,13,21 D.1,4,5,9,14

OpenStudy (jdcap):

A. with r=3

OpenStudy (jdcap):

Since 18/6=54/18=162/54=3